OpenStudy (anonymous):

The product of 9 and a number is 12 less than three times that number. What is the number?

OpenStudy (anonymous):

9+x=3x-12

OpenStudy (anonymous):

9(12)<3X

OpenStudy (anonymous):

I am still completely confused.

OpenStudy (anonymous):

@katti @pages2000

OpenStudy (anonymous):

9x=3x-12 9x+12=3x 12=-6x -2=x